Last year, an online store sold a large number of computers. 55% of the computers were made by company F, 30% were made by company G and 15% were made by company H. A random sample of 3 customers who each bought a computer from this store is chosen.
- (a)Find the probability that the 3 customers bought computers all made by different companies.[1]
- (b)A random sample of 12 customers who each bought a computer from this store is chosen. Find the probability that fewer than 10 of these customers bought a computer made by company F.[3]
- (c)A random sample of 140 customers who each bought a computer from this store is chosen. Use a suitable approximation to find the probability that more than 24 of these customers bought a computer made by company H.[5]
Mark scheme
Marking points from the official mark scheme (Feb/Mar 2025). The mark or mark type is shown in brackets.
- [1] [0.55×0.3×0.15×3!=]0.1485, 297/2000
- Accept 148500/1000000 condone 0.149.
- [1] One term of the form 12Cx (p)^x (1-p)^(12-x), 0 < p < 1, x ≠ 0 or 12.
- [1] Correct unsimplified expression, no terms omitted leading to final answer. Condone omission of last bracket '}' only.
- [1] 0.9575 < p ≤ 0.958.
- [Method 1 Answer] 1 - {12C10 0.55^10 0.45^2 + 12C11 0.55^11 0.45 + 0.55^12} = 1 - (0.0338529 + 0.0075229 + 0.0007662) = 0.957858 = 0.958
- [Method 2 Answer] P(0,1,...,9) = 0.45^12 + 12C1 0.55^1 0.45^11 + ... + 12C9 0.55^9 0.45^3 = 0.958
- [1] 21 and 17.85 (or 177/20) seen, allow unsimplified. May be in standardisation formula. (σ = √17.85, 4.224926...to at least 4SF implies correct variance). Withhold mark if variance clearly identified as standard deviation, condone N(21, √17.85 ) if standardisation formula correct or variance/standard deviation correctly stated as well.
- [1] Substituting their μ and their σ into the ± standardisation formula (any number for 24.5), allow σ² or √σ.
- [1] Use continuity correction 23.5 or 24.5 in their standardisation formula. Note: If no working ±3.5/√17.85 or ±3.5/4.225 seen gains M2 BOD.
- [1] Appropriate area Φ, from final process, must be a probability. May be implied by a sketch of the required probability area. Note: correct final answer implies this M1. Expect final answer < 0.5.
- [1] Final answer AWRT.
- [Answer] Mean = 140 × 0.15 = 21. Variance = 140 × 0.15 × 0.85 = 17.85. P(X > 24) ≈ P(Z > (24.5 - 21)/√17.85) = P(Z > 3.5/4.2249) = P(Z > 0.8284) = 1 - Φ(0.8284) = 1 - 0.7961 = 0.2039. Rounded to 3 significant figures, P(X > 24) = 0.204.
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →